Atomos Test Generator Free PYQs ExamsMy practice Log in / Sign up free

Probability

JEE Main 2026 (Online) 22nd January Morning ShiftMCQEasy📈 Topic in 36% papers+4−1
Ques. 1JEE Main 2026, 22 January (Morning Shift)

Two distinct numbers a and b are selected at random from 1, 2, 3, , 50. The probability, that their product ab is divisible by 3, is

  1. 2721225
  2. 5611225
  3. 6641225
  4. 825

Answer

(3) 6641225

Solution

The total number of ways is 50C2 = 50 492 = 1225.

The multiples of 3 from 1 to 50 are 3, 6, , 48, i.e. 16 numbers, so there are 50 - 16 = 34 non-multiples.

ab is not divisible by 3 only when both numbers are non-multiples, which happens in 34C2 = 34 332 = 561 ways.

P(ab divisible by 3) = 1 - 5611225 = 6641225

Answer: (3)

Keep practising — Probability

Ques. 2JEE Main 2025, 4 April (Morning Shift)

The probability of forming a 12 persons committee from 4 engineers, 2 doctors and 10 professors containing at least 3 engineers and at least 1 doctor, is

  1. 129182
  2. 1726
  3. 1926
  4. 103182
Ques. 3JEE Main 2025, 2 April (Morning Shift)

Three distinct numbers are selected randomly from the set \1, 2, 3, , 40\. If the probability, that the selected numbers are in an increasing G.P., is mn, (m, n) = 1, then m + n is equal to .

Ques. 4JEE Main 2025, 28 January (Morning Shift)

Two numbers k1 and k2 are randomly chosen from the set of natural numbers. Then, the probability that the value of ik1 + ik2, (i = -1) is non-zero, equals

  1. 34
  2. 12
  3. 14
  4. 23
Ques. 5JEE Main 2025, 28 January (Evening Shift)

Let S be the set of all the words that can be formed by arranging all the letters of the word GARDEN. From the set S, one word is selected at random. The probability that the selected word will NOT have vowels in alphabetical order is:

  1. 14
  2. 12
  3. 13
  4. 23
Ques. 6JEE Main 2025, 23 January (Morning Shift)

One die has two faces marked 1 , two faces marked 2 , one face marked 3 and one face marked 4 . Another die has one face marked 1 , two faces marked 2 , two faces marked 3 and one face marked 4. The probability of getting the sum of numbers to be 4 or 5 , when both the dice are thrown together, is

  1. 23
  2. 35
  3. 49
  4. 12
Ques. 7JEE Main 2025, 23 January (Evening Shift)

A board has 16 squares as shown in the figure : [figure] Out of these 16 squares, two squares are chosen at random. The probability that they have no side in common is :

  1. 35
  2. 45
  3. 2330
  4. 710
Ques. 8JEE Main 2024, 9 April (Morning Shift)

Let a, b and c denote the outcome of three independent rolls of a fair tetrahedral die, whose four faces are marked 1, 2, 3, 4. If the probability that ax2 + bx + c = 0 has all real roots is mn, (m, n) = 1, then m + n is equal to .

Ques. 9JEE Main 2024, 9 April (Evening Shift)

If an unbiased dice is rolled thrice, then the probability of getting a greater number in the ith roll than the number obtained in the (i - 1)th roll, i = 2, 3, is equal to

  1. 5/54
  2. 2/54
  3. 1/54
  4. 3/54
Ques. 10JEE Main 2024, 8 April (Morning Shift)

Let the sum of two positive integers be 24. If the probability, that their product is not less than 34 times their greatest possible product, is mn, where (m, n) = 1, then n - m equals

  1. 10
  2. 11
  3. 9
  4. 8
Ques. 11JEE Main 2024, 6 April (Evening Shift)

If three letters can be posted to any one of the 5 different addresses, then the probability that the three letters are posted to exactly two addresses is :

  1. 1825
  2. 1225
  3. 625
  4. 425

All Probability PYQs →

PDFDownload Probability PYQs as PDF Free, with answer key and a QR code to every solution — print it or study offline⬇
All JEE Main Mathematics chapters